OpenStudy (staceyg):

The sum of the measures of angle X and angle Y is 90. If the measure of angle X is 30 less than twice the measure of angle Y, what is the measure of angle X?

OpenStudy (anonymous):

make the equation for this

OpenStudy (anonymous):

x + y = 90 x = 2y - 30

OpenStudy (campbell_st):

you you have angle y then x = 2y - 30 so adding the complementary angles y + 2y - 30 = 90 collect like terms and solve for y

OpenStudy (campbell_st):

when you have y, you can find x

OpenStudy (campbell_st):

why just give the answers...

OpenStudy (staceyg):

I got y=20 when solved your equation but then that would mean x=10.

OpenStudy (campbell_st):

well when you collect like terms you get 3y - 30 = 90 add 30 to both sides gives 3y = 120 divide by 3 to get y

OpenStudy (staceyg):

Oh right! I did that wrong. Thanks for the help!

OpenStudy (campbell_st):

glad to help... good luck